A new recreational area in a small town will take 20 full-time (40 hours per week) workers 3 weeks to complete. In addition, the cost of materials is expected to be $75,000.

a. If the labor market is perfectly competitive and the market wage is $20, what are the total opportunity costs of taking on this project? SHOW YOUR WORK AND EXPLAIN?

b. If instead the 20 individuals hired were unemployed and valued leisure at $10 per hour, would the opportunity cost of the project change? SHOW YOUR WORK AND EXPLAIN DIFFERENCE BETWEEN (a) and (b)
Answer #1

Solution

Number of workers = 20

Working hours = 40 hours per week

Duration = 3 weeks

Cost of material = $ 75000

(a) When the labor market is perfectly competitive

Market wage = $20

Cost of Labor = No of Workers * Working hours * Duration = 20*40*3 = 2400

Total Labor Cost = 20 *2400 = $48000

The total opportunity costs of taking on this project are equal to

= Total labor cost + material cost

= $(48000 +75000) = $123000

(b)

Yes, there will be a change in the opportunity cost when 20 individuals hired were unemployed and valued leisure at $10 per hour. These unemployed individuals will be working as a laborer will value their time for working less than $20 an hour, which in turn leads to lowering of the opportunity cost of the project.